### v1

# Black Big Boxes: Do Language Models Hide a Theory of Adjective Order?

Abstract
In English and other languages, multiple adjectives in a complex noun phrase show intricate ordering patterns that have been a target of much linguistic theory. These patterns offer an opportunity to assess the ability of language models (LMs) to learn subtle rules of language involving factors that cross the traditional divisions of syntax, semantics, and pragmatics. We review existing hypotheses designed to explain Adjective Order Preferences (AOPs) in humans and develop a setup to study AOPs in LMs: we present a reusable corpus of adjective pairs and define AOP measures for LMs. With these tools, we study a series of LMs across intermediate checkpoints during training. We find that all models’ predictions are much closer to human AOPs than predictions generated by factors identified in theoretical linguistics. At the same time, we demonstrate that the observed AOPs in LMs are strongly correlated with the frequency of the adjective pairs in the training data and report limited generalization to unseen combinations. This highlights the difficulty in establishing the link between LM performance and linguistic theory. We therefore conclude with a road map for future studies our results set the stage for, and a discussion of key questions about the nature of knowledge in LMs and their ability to generalize beyond the training sets. [^1] All code and data, including CAP, can be found here: github.com/jumelet/lm-adjorder
1 Introduction
Figure 1: We connect the adjective order preferences (AOP-$\Delta$, §3.1) of language models (here Pythia-12b) to the adjective order frequencies of the corpus they have been trained on (The Pile). We highlight various regions of interest: adjective pairs for which both orders are rare and that require the LM to generalize from other adjective orders; pairs for which one particular order is far more common that can be resolved from frequency alone; and orders with high variance that rely more strongly on context.
Linguistic capabilities of language models (LMs) have been subject to a lot of research and benchmarking (e.g., Marvin and Linzen 2018; Hu et al. 2020; Gauthier et al. 2020). While LMs show impressive fluency in language generation and often score high on linguistic benchmarks, the mechanisms underlying their linguistic proficiency are hard to uncover. To what extent do LMs generalize beyond mere memorisation of training data? Do they acquire abstract linguistic rules and constraints that don’t boil down to basic co-occurrence patterns? How similar are linguistic generalizations in LMs to those that humans have?
These questions have been asked and partially answered for different constructions and grammatical phenomena. The focus of such studies has mostly been on strict and binary grammatical constraints, where a violation results in an ungrammatical sequence.
There has been less attention to how ‘softer’ linguistic constraints are captured by LMs. Such constraints can give rise to slight and defeasible preferences rather than sharp acceptability contrasts. One example is adjective order in noun phrases where multiple adjectives modify one noun, e.g. large wooden box. Changing the order of adjectives does not necessarily lead to unacceptability, but can result in a slight decrease in naturalness (Dyer et al., 2023) (‘>’ marks preference):
  1. large wooden box $>$ wooden large box
Moreover, these preferences are context-dependent: often, a context can be found where preferences switch – for instance, when a contrasting property is introduced (Teodorescu, 2006).
  1. Take a wooden large box, not a plastic one.
  2. $>$   … large wooden box …
These adjective order preferences (AOPs) have been subject to a lot of interest in theoretical linguistics, but proved difficult to capture due to their graded and context-sensitive nature (see Levshina et al. 2023 for arguments in favor of a systematically gradient approach to word order). A number of factors ranging from superficial properties like adjective length to semantic and pragmatic properties to information-theoretic properties of a sentence have been identified as predictors for AOP. But even with all those factors combined together, a large portion of AOPs remains unaccounted for: the best model combining different predictors explains only 71% of human AOPs (Dyer et al., 2023).
Adjective order offers an opportunity to assess the ability of LMs to learn subtle linguistic preferences involving implicit, hard-to-identify factors. This paper is the first study of AOP in LMs. We put together metrics and data, and conduct a series of experiments that explore AOP in LMs, its relation to properties of training data and the extent to which LMs learn more abstract principles behind AOP, generalizing to unseen items. The contributions of our paper are the following:
[…]
  • We show that LMs are impressively good predictors of naturally occurring adjective orders, with AOP prediction accuracy as high as 94.1%. We identify three stages of AOP acquisition in LMs that lead to their performance at the end of the training (§4).
  • We demonstrate that linguistic and cognitive predictors only partly account for LM performance in AOP prediction. We identify the gap between linguistic theory and LM performance in the AOP domain that needs to be explained (§5).
  • To pinpoint the role of training data co-occurrence statistics in LM AOP predictions, we study the Pile (Gao et al., 2020) – a large-scale corpus that Pythia models were trained on. Simple bigram statistics turn out to be a powerful AOP predictor for naturally occurring adjective orders on their own, at 90.3% accuracy (§6). Figure 1 summarizes a close relation between count statistics and LM predictions, questioning whether LM predictions are significantly driven by anything beyond $n$-gram counts.
  • At the same time, we find two ways in which LMs go beyond simple collocational statistics: 1) We test LM predictions for adjectival combinations not seen during training (a challenging setup due to the large size of the Pile; we overcome this obstacle by turning to intermediate checkpoints and corresponding portions of training data). We find some—but limited—generalization to new data (§6); 2) We study the effect of context on AOP. We conclude that the sentence helps LMs make better AOP predictions compared to noun phrase with adjectives taken in isolation or in a set of random contexts, which indicates that AOP mechanisms in LMs go beyond $n$-gram counts in training data (§7).
2 Background
2.1 Adjective Order Theory
The mechanisms that drive adjective order have been a topic of research for decades (see Dyer et al. 2023 for overview). Theories have been proposed as far back as the 19th century (Sweet, 1898), and refinements to the theory are being made to this day. Early theories proposed lexical hierarchies that described adjective order based on abstract classes such as dimension or color (Dixon, 1982; Sproat and Shih, 1991; Cinque, 1996; Scott, 2002). Adjectives were expected to be ordered based on the class they belong to: e.g. dimension goes in front of color, which gives rise to the order of a phrase like the big red house.
While these hierarchies are powerful predictors of adjective order, they are not really explanatory: the question remains how such hierarchies are formed by the underlying properties of adjectives. Furthermore, corpus studies have shown that many counterexamples against various versions of these hierarchies can be found (Truswell, 2009), demonstrating the need for a more fine-grained analysis (Trotzke and Wittenberg, 2019).
One line of work looks for semantic driving forces of AOP, placing emphasis on adjectival properties like intersectivity and subsectivity (Truswell, 2009) or identifying conditions under which adjective order constraints can be lifted, such as focus (Vendler, 1968; Teodorescu, 2006).
Other approaches take a broader perspective, leveraging insights from quantitative linguistics, distributional semantics and information theory, and connecting this to psycholinguistic theories of human language processing. Early analyses focus on properties such as length—shorter adjectives tend to precede longer ones (Behaghel, 1930)—and frequency: more frequent adjectives tend to appear earlier (Martin, 1969; Ney, 1981; Wulff, 2003). This follows from a broader theory of language production, according to which more accessible and familiar material is produced first (Bock, 1982; Ferreira and Dell, 2000) and is comprehended faster (Arnon and Snider, 2010). Collocational patterns of adjectives and nouns have been shown to be highly predictive of adjective order: adjectives most closely connected to the noun in term of co-occurrence appear closer to it (Sweet, 1898; Byrne, 1979; Lapata et al., 1999). This intuition has been operationalized through pointwise mutual information (PMI) (Hahn et al., 2018; Futrell et al., 2020a), which aligns with more general theories of word order arguing that distance between words that closely predict each other tends to be small (Gibson, 1998; Futrell et al., 2020b).
Predictors themselves can be complex, based on collocational or semantic factors or combinations thereof: modification strength of adjectives, for instance, is a predictor based on adjectival compositional potential (Vecchi et al., 2013, 2017); other predictors proposed in the literature include information gain that an adjective provides for a noun (Dyer et al., 2021), and subjectivity of adjectives, with more subjective adjectives occurring first (Hill, 2012; Scontras et al., 2017, 2019; Franke et al., 2019). Although the majority of these approaches focus on English, various cross-lingual accounts have demonstrated their universal nature: adjective order tends to be driven by the same principles across languages (Wulff and Gries, 2015; Leung et al., 2020; Scontras, 2023; Dyer et al., 2023).
Finally, various papers highlight the need for a multi-factorial account of adjective order: it is unlikely adjective order is driven by a single factor and therefore it’s the result of various competing pressures (Wulff, 2003; Trotzke and Wittenberg, 2019; Dyer et al., 2023). Relatedly, a recent position paper (Levshina et al., 2023) argues for a fundamentally gradient theory of word order that would be able to capture its hybrid and multi-faceted nature.
In our work, we compare LM predictions to various adjective order theories. Additionally, we examine the utility of LMs for investigating the role that context plays on word order: since almost all theories focus on adjectives in isolation, LMs may provide a fruitful starting point for shaping a contextual account of adjective order.
2.2 Word Order in Language Models
With the increasing capabilities of language models, their linguistic abilities have been under much scrutiny in recent years. A wide range of procedures have been introduced to uncover their understanding of grammaticality and notions of linguistic structure (Linzen and Baroni, 2021). Grammaticality is often assessed using the LM’s probability predictions of minimal pairs of grammatical and ungrammatical sentences (Marvin and Linzen, 2018; Warstadt et al., 2020; Gauthier et al., 2020); phenomena that have been examined this way include number agreement (Linzen et al., 2016; Gulordava et al., 2018), negative polarity items (Jumelet and Hupkes, 2018; Bylinina and Tikhonov, 2022), and filler-gap dependencies (Wilcox et al., 2018; Suijkerbuijk et al., 2023). To our knowledge, no prior work has focused on adjective order in particular: the closest related phenomenon to this would be Misra and Mahowald (2024)’s analysis of adjective-numeral constructions in LMs.
The importance of word order in LMs has been a topic of debate, with various works claiming that downstream performance is not affected by scrambled inputs (Malkin et al., 2021; Sinha et al., 2021), although it has been shown that LMs are able to retain a notion of word order through their positional embeddings (Abdou et al., 2022). It has been argued that LMs acquire an abstract notion of word order that goes beyond mere $n$-gram co-occurrence statistics (Futrell and Levy, 2019; Kuribayashi et al., 2020; Merrill et al., 2024), a claim that we in this paper assess for large-scale LMs in the context of adjective order. Finally, numerous works have investigated the trade-off between memorization and generalization in LMs: it has been shown that larger LMs are able to memorize entire passages from the training data (Biderman et al., 2023a; Lesci et al., 2024; Prashanth et al., 2024), but generalization patterns for grammatical phenomena have also been shown to follow human-like generalization (Dankers et al., 2021; Hupkes et al., 2023; Alhama et al., 2023).
3 Methods
3.1 Measuring Word Order Preference
We measure a language model’s AOP by comparing its log probability of the natural order (the order in which a pair of adjectives appears in the sentence from our corpus) with a swapped order. In our experiments we measure AOP in two settings, either by considering the phrase in isolation [^2] We compute these isolated probabilities by prepending the noun phrase with ‘The’., or by taking the context of the phrase into account. For a double adjective phrase $\textsc{a}_{1}\textsc{a}_{2}\textsc{n}$ that is extracted from a corpus with context c, we compute a model’s preferred order as follows:
[…]
All probabilities are in log space. The AOP-$\Delta$ metric represents the magnitude of a models’ preference. In our experiments we also consider an accuracy metric based on AOP-$\Delta$, expressed as the number of items for which AOP-$\Delta$ is positive:
$\textup{AOP}\textit{-\%}(\mathcal{C})=\frac{\left\{\phi\in\mathcal{C}\mid\textup{AOP-}\Delta(\phi)>0\right\}}{|\mathcal{C}|}$ (3)
Where $\mathcal{C}$ is an evaluation corpus of double adjective phrases, as described in the following section. Misra and Mahowald (2024) consider a similar metric for measuring properties of word order preferences in LMs, based on corrupted word orders.
[…]
Procedure
As a starting point, we use the BabyLM 10M corpus (Warstadt et al., 2023), which consists of a mix of various English language sources including Wikipedia articles, books, and transcribed dialogue data. From this corpus we make a pre-selection of sentences containing multiple adjectives by filtering all sentences containing at least two adjectives, based on a list of 11,000 adjectives released by Futrell et al. (2020a). Next, we generate dependency parse trees for each sentence using spaCy and select all sentences in which a noun contains two amod relations to an adj that is part of the adjective list. This procedure results in 2678 double adjective sentences. We provide a sample of CAP sentences in Appendix A.1.
From these sentences we also generated corrupted sentences in which the adjective order is swapped. In case the first adjective is preceded by an indefinite article we ensure the article is changed accordingly to satisfy morpho-phonological constraints. Here is an example of an extracted phrase and its swapped counterpart:
  1. $+$
    Peter received a full athletic scholarship [..]
  2. --
    Peter received an athletic full scholarship [..]
3.3 Models
For our experiments we make use of the Pythia suite of language models (Biderman et al., 2023b). The models are released in increasing sizes (from 70M to 12B parameters) and all trained on The Pile corpus (Gao et al., 2020). Importantly, all intermediate checkpoints are available for these models in logarithmic intervals, which allows us to investigate the learning dynamics at a fine-grained level.
Figure 2: A–B: AOP-and AOP-$\Delta$ scores for Pythia models of increasing size. C–D: AOP-and AOP-$\Delta$ scores for Pythia-1.4b during training. We highlight the three learning phases: 1) initialization, 2) acquisition, and 3) consolidation.
4 Experiment 1: AOP in LLMs
In this first set of experiments we evaluate a series of pretrained LLMs on CAP using the AOP-$\Delta$ and AOP-metrics of §3.1.
Model Size
For the eight Pythia models of increasing size we plot the AOP-performance in Figure 2A, and the AOP-$\Delta$ performance in Figure 2B. It can be seen that performance improves with model size: contextual AOP-performance increases from 89.6% for the 70m model to 94.1% for the 12b model. AOP-does not increase consistently with size, however, from the 1b model onward it reaches a plateau. Interestingly, AOP-$\Delta$ does increase consistently with size: this shows that larger models become more certain about their AOP judgment.
Learning Dynamics
To investigate how AOP develops during training, we compute scores for the 1.4b model on all its intermediate checkpoints in Figure 2C and 2D. From these plots we can identify three distinct learning phases: 1) an initialization phase in which AOP judgments have not been formed yet, 2) an acquisition phase in which AOP judgments are rapidly being formed, and 3) a consolidation phase in which AOP judgments remain stable and only the role of context is being reinforced. A similar three-phase learning dynamic has been observed by Van Der Wal et al. (2022) and Chen et al. (2023). Note that the consolidation phase is reached relatively quickly: after around 2000 batches, which is ~1.4% of the total model training. This demonstrates that adjective order is a relatively low-level linguistic feature that is acquired early during training. Furthermore, it highlights the importance of investigating learning dynamics at a logarithmic scale: with linear checkpoints we would not obtain this fine-grained level of precision.
Figure 3: Average token probabilities for the original and swapped adjective orders on Pythia-12b, without and with sentence context (A–B), as well as the token-level differences (C) that correspond to the difference between the curves in (A) and (B).
Localizing AOP
In our experiments so far we have measured AOP based on the probabilities of the full adjective-noun triplet. Preference order can arise at different locations of the phrase. The most natural location for this is the probability of the second adjective after having seen the first adjective, but context could also play a role for preferring which adjective comes at the first position, and a noun could be more probable after having seen two adjectives in their natural order.
In Figure 3 we plot the average token probabilities on CAP for Pythia-12b, in isolation and contextual. We observe that, as expected, the largest probability difference occurs at the position of the second adjective: the model is highly surprised by encountering the second swapped adjective after having seen the first. The noun probability is strongly affected by swapping adjective order as well: a noun is more likely to follow adjectives in natural order. Furthermore, we can see that context plays a role for the probability of the first adjective already: whereas in the isolated case both adjectives in first position receive equal probability, order preference already manifests here when the model has access to context.
Conclusion
The Pythia LMs acquire a strong sense of adjective order that is learned early in training at around 1.5% of the total amount of training. We showed that sentence context has a positive effect on AOP, bumping up AOP-from 91% to around 94%. By localizing the AOP scores on the token-level we show how these order preferences manifest at different points in a phrase. In the next experiments we will investigate these results in more detail, connecting LM AOP to cognitive predictors and training data statistics.
5 Experiment 2: Cognitive Predictors
[…]
Metric AOP-$\rho(\textup{AOP-}\Delta(\bullet))$ $\rho(\textup{AOP-}\Delta(\bullet|c))$
AOP-$\Delta(\bullet)$ 91.6 0.71
[…]
PMI(a;n) 62.4 0.27 0.20
Subj(a) 69.3 0.07 0.08
[…]
%$\%(\textsc{a}_{1}\textsc{a}_{2})$ 90.3 0.58 0.53
%$\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})$ 87.6 0.54 0.53
Table 1: Predictive accuracy of adjective order on CAP for various metrics: AOP-$\Delta$ for Pythia-12b, 3 cognitive predictors (§5), and relative n-gram counts from The Pile (§6). The highest accuracy for each group is bolded. For each metric we provide the Spearman correlations with respect to the AOP-$\Delta$ scores of Pythia-12b.
As discussed in §2.1, a range of cognitively plausible predictors have been proposed to explain adjective order. In the previous section we have shown that LMs are highly capable of predicting adjective order, and in this section we investigate how their predictions are related to various theories on adjective order. We evaluate the performance of three cognitive predictors—Length, PMI, and Subjectivity— and correlate them against AOP-$\Delta$ scores across training.
5.1 Cognitive Predictors
For each predictor we compute its score for a particular adjective pair. An adjective order predictor $f$ returns a graded score for adjective order, akin to our formulation of AOP-$\Delta$.
Length
For Length—which hypothesizes that the first adjective is the shortest—the score for adjectives a1 and a2 is expressed as
$f(\textsc{a}_{1},\textsc{a}_{2})=|\textsc{a}_{2}|-|\textsc{a}_{1}|$
We also considered expressing length as the number of tokenizer subwords, since this is the only notion of length LMs have access to, but found that this led to similar results as string length.
PMI
We follow the procedure of Futrell et al. (2020a) for computing PMI, which is based on the distribution over adjective relations in dependency parse trees. It is defined as
$\textup{PMI}(\textsc{a};\textsc{n})=P(\textsc{a},\textsc{n})/(P(\textsc{a})\cdot P(\textsc{n}))$
$P(\textsc{a},\textsc{n})$ is computed as the normalized number of occurrences of adjective a and noun n in amod relation. $P(\textsc{a})$ and $P(\textsc{n})$ are obtained by marginalizing over this joint distribution. PMI hypothesizes that adjectives closer to the noun have a higher PMI:
$f(\textsc{a}_{1},\textsc{a}_{2},\textsc{n})=\textup{PMI}(\textsc{a}_{2};\textsc{n})-\textup{PMI}(\textsc{a}_{1};\textsc{n})$
Subjectivity
We leverage the 450 subjectivity ratings provided by Dyer et al. (2023), to compute subjectivity scores for the subset of adjective pairs that contain a rating. This procedure retains 37.5% of the adjective pairs in CAP (1005 items). Subjectivity as predictor hypothesizes that more subjective adjectives occur first (Scontras et al., 2017):
$f(\textsc{a}_{1},\textsc{a}_{2})=\textup{Subj}(\textsc{a}_{1})-\textup{Subj}(\textsc{a}_{2})$
5.2 Predictive Accuracy
We present the predictive accuracy of each predictor in Table 1. Subjectivity is the strongest predictor with 69.3% accuracy, followed by PMI (62.4%) and Length (49.6%). While Subjectivity reaches the highest accuracy, its correlation with the LM’s AOP-$\Delta$ scores is surprisingly the lowest of the three. We plot these correlations over time in Appendix A.2: all three correlations develop in the acquisition phase and remain stable in the consolidation phase.
We also compute the overlap between these three predictors on the 1005 items for which we have subjectivity scores:
[Uncaptioned image][Uncaptioned image]
The first Venn diagram makes clear that the cognitive predictors have low agreement: only 213 orders are predicted correctly by all three. The broad coverage (93.1%) of the union, however, shows that most of the CAP items are explained by at least one of these predictors. In the four diagrams on the right we show the overlap of the predictors with the AOP-$\Delta$ scores of Pythia-12b. All three predictors and their intersection are almost fully subsumed by the AOP-$\Delta$ predictions. This demonstrates that AOP-$\Delta$’s unique contribution goes significantly above that of the ensemble of the cognitive factors, raising the question whether LMs ‘hide’ a meaningful linguistic theory that extends beyond established factors.
6 Experiment 3: Is AOP Driven by Data Statistics?
Figure 4: Correlations during training of LM probabilities for single adjectives, adjective pairs, and adjective-noun triplets with respect to their frequency in The Pile.
In order to truly understand what LMs base their adjective order preferences on, we need to get a better insight into how double adjective constructions are distributed in the training data. The way that contemporary LMs are trained nowadays makes this challenging: many models are released without open access to their training data, and the scale of training corpora that are available poses a challenge for targeted analysis.
6.1 WIMBD
To overcome this, we make use of the WIMBD API (What’s In My Big Data, Elazar et al., 2023), which provides a fully indexed search engine over various large-scale corpora such as The Pile, on which the Pythia models have been trained. WIMBD makes it possible to retrieve the exact $n$-gram statistics that LMs have been trained on, and as such provides a detailed insight into the collocational factors that play a direct role in shaping AOP.
We retrieve the uni-, bi- and trigram counts for all the adjective-noun triplets in CAP for The Pile, as well as for the swapped orders. To provide an analogue to AOP-$\Delta$, we express the n-gram counts as log count differences. We express these relative n-gram counts as
%$\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})=\log\#(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})-\log\#(\textsc{a}_{2}\textsc{a}_{1}\textsc{n})$
with similar formulations for adjective bigrams (%($\textsc{a}_{1}\textsc{a}_{2}$)) and unigrams (%($\textsc{a}_{1}$)).
Predictive Accuracy
We present the predictive accuracy of these relative n-gram occurrences in Table 1. %(a1) predicts adjective order with 60.5%, which provides weak evidence that more frequent adjectives tend to occur first (Martin, 1969; Trotzke and Wittenberg, 2019). %(a1a2) is the strongest predictor of adjective order—except for the LMs—at 90.3%. This shows that a large majority of the items in CAP can be resolved based on bigram statistics alone, which is leveraged by the LMs during training. However, n-gram collocations alone do not fully explain LM performance: a large amount of trigrams are not part of the Pile and performance for %$\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})$ drops to 87.6%.
Correlations
To investigate how n-gram statistics play a role in the formation of AOP-$\Delta$ during training, we compute the Spearman correlation between various log-transformed n-gram counts and a LM’s AOP-$\Delta$ on intermediate checkpoints. We plot this in Figure 4, for contextual and isolated AOP-$\Delta$. The trends we observe here align closely with the three learning phases that we outlined in §4. The initialization phase (1) is marked by a rapid acquisition of unigram and bigram statistics, that remain unaffected by context at this point. Then, at the start of the acquisition phase (2), in which AOP performance rapidly improves, the model starts to incorporate linguistic context in its uni- and bigram predictions. Finally, in the consolidation phase (3), trigram predictions start being affected by context as well, whereas unigram and bigram predictions remain relatively stable. The fully trained model strongly reflects these correlations to data statistics, which explains the strong relation we observed in Figure 1 between bigram frequencies and AOP-$\Delta$.
6.2 Generalizing to Unseen Adjective Orders
Figure 5: The contextual AOP-performance for Pythia-70m and 1.4b across training, split out for items that have been seen 0, 1, 2 to 10, and more than 10 times at each specific checkpoint.
Having access to the n-gram distribution a model was trained on allows us to investigate its performance on unseen adjective orders: is the model able to generalize from an abstract signal of adjective order to novel orders? However, due to the enormous size of The Pile (300B tokens) it turns out that number of unseen adjective pairs in CAP is too small to draw significant conclusions from (only 4 out 2678 pairs). This leaves us with two options: either we collect an additional sample of rare adjective pairs that are not part of the Pile, or we investigate the model at intermediate checkpoints where a greater amount of adjective pairs has not been encountered yet. An issue with collecting highly rare pairs is that their canonical order will be much harder to determine, for which we would require extensive human judgments on the most natural order. We therefore go for the second option, and set up a procedure to measure the data statistics during training.
Intermediate $n$-gram Counts
Since WIMBD only provides $n$-gram statistics for The Pile as a whole, we need to collect the intermediate statistics ourselves. Using the Pythia batch viewer (Biderman et al., 2023b) this becomes possible: our implementation is able to process the first 10% of The Pile (up to batch 14,000; 29.4B tokens) in around 3 hours. We collect the bi-gram counts of all adjective pairs in CAP on the batch level, allowing us to determine exactly when and how often a specific adjective pair has been seen by the model.
Unseen Adjective Pairs
Based on these intermediate $n$-gram counts, we compute the AOP-scores for various splits based on the number of times an adjective pair has been encountered. We focus on the cases where the swapped order has never been encountered: this allows us to isolate the moment a model encounters an adjective combination for the first time, and measure its ‘zero-shot’ performance on cases that are not already skewed towards to swapped order. For each model checkpoint we collect the $n$-gram counts at that particular point in training, and split the CAP items based on the number of times an a1a2 bi-gram has been seen. We provide the size of these splits across training in Appendix A.3.
We present the results for Pythia-70m and 1.4b in Figure 5. The performance on unseen pairs is remarkably high, reaching around 85% at the highest point for both models. Interestingly, the gap between unseen pairs and pairs seen once is far greater for the 1.4b model than the small 70m model. This shows that this larger model is able to acquire adjective order preference faster: a single exposition to an order immediately leads to an increase in AOP-$\Delta$. We provide a more detailed item-level analysis for these items in Appendix A.4: while it is clear on the split level that a single occurrence improves AOP-%, we do not observe such a clear effect on the item level. Interestingly, as the number of unseen adjective pairs decreases, the performance on unseen pairs decreases as well. This shows that the predicting the order for the set of unseen pairs around batch 2,000 (6% of CAP) could be better inferred from a general signal of adjective order than the set at batch 14,000 (1.3% of CAP).
Conclusions
The question we set out to answer in this set of experiments is whether AOPs are driven by data statistics. We approached this question both in relation to the $n$-gram distribution of the Pile and to the shifting $n$-gram distribution across training. On the one hand, AOP is strongly driven by data statistics: we find a strong correlation between AOP-$\Delta$ and count-based AOP scores, which is also visualized in Figure 1. However, the relatively strong performance on the set of unseen adjective pairs during training indicates that LMs are able to infer order from a more general signal for adjective pairs it has not encountered before.
7 Experiment 4: The Role of Context
Figure 6: A: Item-level AOP-$\Delta$ scores with and without sentence-context for Pythia-12b. We provide a sample of sentences from each quadrant in Appendix A.1. B: The relative impact ($c$-$\Delta$, Eq. 5) of an item’s true context compared to an expectation over random contexts (Eq. 4).
In §4 we established that LMs have a thorough grasp of adjective order that manifests itself early during training. We showed that having access to sentence context boosts performance, and we now investigate the role of context in more detail.
Contextual Improvements
To gain a better insight how much items are impacted by context, we investigate AOP-$\Delta$ scores on an item-level with and without context. In Figure 6A, we plot the AOP-$\Delta$ scores in isolation against the contextual scores, and mark the four quadrants of items that result in a positive or negative score after adding context. In the majority of cases (89.4%), AOP-$\Delta$ is positive both with and without context: the model is able to determine adjective order based on the adjective-noun triplet alone. In 4.8% of the cases, however, adding context flips a negative AOP-$\Delta$ score to a positive one. Adding context has a boosting effect in general: in 65.1% of the cases, AOP-$\Delta$ is increased by adding context. We provide an overview of how these ratios are affected by model size in Appendix A.5: the number of cases where context improves AOP-is actually larger for small models, but so is the number of cases where context worsens it.
Random Contexts
The fact that context improves AOP-$\Delta$ raises the question what information in the context is responsible for this improvement. Does context contain relevant semantic information that plays a meaningful role for determining adjective order, or is any linguistic signal sufficient for this? To test this, we compute an expectation over AOP-$\Delta$ scores with random contexts, in which an adjective pair is preceded by a randomly sampled context from another adjective pair in CAP:
[…]
We compute the expectation over a random sample of 500 contexts, taken from CAP ($\mathcal{C}$). To express the relative impact of adding context, we regress out a model’s isolated AOP-$\Delta$ score by subtracting it from the contextual scores, which we call $c\textup{-}\Delta$:
$c\textup{-}\Delta(\bullet|c)=\textup{AOP-}\Delta(\bullet|c)-\textup{AOP-}\Delta(\bullet)$ (5)
By computing this score for both the original context and for the expectation over random contexts we can quantify how much more important the original context is for AOP-$\Delta$.
We plot the results for this experiment in Figure 6B. The true context improves isolated AOP-$\Delta$ in 65.1% of the cases, whereas random contexts lead to an improvement for only 39.6%. Furthermore, for 33.2% the true context leads to an improvement whereas random contexts don’t: this subset can serve as a valuable starting point for investigating what properties of the context are important for adjective order.
Conclusions
AOP in Pythia is context-sensitive. Providing the model with the sentence where the adjective combination appears increases the model’s preference towards the target adjective order – both compared to the noun phrase in isolation and to the noun phrase in a random context. This result points in the direction of a more linguistically interesting mechanism for AOP than reliance on $n$-gram counts in training data alone. In particular, it shows a non-trivial interaction between the noun phrase and bigger context that turns out relevant for adjective placement. Establishing this opens up a possibility of studying the forms and types of this contextual interaction – a potential novel contribution to linguistic theories of adjective order.
We leave a systematic discovery of types of AOP context-dependence open for future work, and focus on a small-scale qualitative study for now. An inspection of context-sensitive cases from CAP reveals at least two ways in which context matters for AOP:
1. Structural parallelism
The ‘canonical’ word order can be overridden when context contains a parallel noun phrase with adjectives in non-canonical order (the example shows contextual AOP according to Pythia-1.4b):
  • a)
    The green big cars AOP-$\Delta$ 4.2
  • b)
    The story is not about small red houses, it’s about big green cars AOP-$\Delta$ 6.63
2. Property saliency
The property salient in the context brings the adjective describing this property to the left edge of the noun phrase. If the sentence is about color, color adjective will come first; if the sentence is about size, it will bring size adjectives to the left (again, orders preferred by Pythia-1.4b):
  • a)
    They asked my favourite color of houses and I said I liked red small houses AOP-$\Delta$ 0.32
  • b)
    They asked my favourite size of houses and I said I liked small red houses AOP-$\Delta$ 1.61
These observations, of course, barely scratch the surface of possible contextual AOP interactions, but exemplify the power of LMs in incorporating contextual information into word order judgments.
8 Discussion
Abstract Notions of Adjective Order
In this paper, we looked for AOP generalization in LMs. We were interested in whether abstract driving forces are at play in LM predictions beyond simple $n$-gram counts in training data. From our experiments we conclude that LMs show limited AOP generalization. Here, we want to highlight an important question that is rarely brought up in linguistic generalization discussions in NLP: How general and abstract is human linguistic behavior?
Linguistic literature suggests that human language production, including word order preferences, is guided by a mix of memorization and abstract principles (O’Donnell, 2015). Speakers tend to recycle more routinized and therefore more accessible combinations (MacDonald, 2013), guided by sequence frequencies in linguistic input that, in turn, form sequential associations that lead to production automatizing (Bybee, 2010; Diessel, 2019). This is compatible with a picture of language acquisition where knowledge of grammar is formed based on encountered exemplars of usage, which are stored in memory (Bybee, 2010) and serve as basis of generalization (Goldberg, 2006), but can also serve as patterns for routinized language production. So, human linguistic knowledge and language production is not totally abstract and is not based exclusively on generalization. This has been discussed across linguistic phenomena and has been suggested for word order (MacDonald, 2013; Levshina et al., 2023), although not for adjective order specifically.
When investigating generalization in LMs it is therefore important to make a distinction between the degree to which they create generalizations for a particular phenomenon, and how similar this generalization is to human behavior. In principle, one could even find cases of LM overgeneralization, if we are interested in how human-like LMs are linguistically. Importantly, in order to answer the question of human-like generalization, we would need data on the trade-off in memorization and generalization in human linguistic behavior. Such data does not exist yet for adjective order preferences.
LMs and Linguistic Theory
Despite the impressive performance of LMs in recent years, their impact on theoretical linguistics has remained minimal (Baroni, 2022). There are various reasons for this. On the one hand, the cognitive plausibility of current LMs remains questionable, both in terms of the scale at which they are trained (Huebner et al., 2021; Warstadt et al., 2023) and their inductive biases that depend strongly on model architecture and are not reflective of human language processing (Wilson and Frank, 2023; Oh et al., 2024). On the other hand, their black box nature makes it highly challenging to extract linguistic patterns that can be tested against existing linguistic theories. While advancements in interpretability have made this latter point less problematic, the number of works explicitly connecting LM behavior to human language processing remains few and limited to simple syntactic phenomena (e.g., Lakretz et al., 2021).
In this paper, we explore the possibility that LMs may ‘hide’ a theory of adjective order, by encoding features that are not covered by linguistic theories on adjective order. We show that LMs far exceed the predictive power of prior cognitive predictors, which raises the question whether this increase stems from meaningful interpretable features. While a lot of the LM predictions can be explained by collocational patterns—and those statistics alone do not explain the distribution of adjective orders—we do find some generalization to unseen adjective combinations and a linguistically meaningful pattern in the way that context boosts AOP. We hope our work can serve as an inspiration for using LMs as tools for measuring context-dependence, which is hard to quantify with traditional quantitative methods.
Corpus Interventions
Our analysis of AOP abstractness is constrained by the fixed training corpus the models are trained on. More detailed insights into this question can be obtained by a intervention-based approach, in which we filter out specific constructions from the training data (Jumelet et al., 2021; Misra and Mahowald, 2024; Patil et al., 2024). An interesting experiment in this direction would be to remove all sentences containing double adjectives from the training corpus, and measure whether the remaining distribution of adjectives contains a sufficient signal to acquire adjective order constraints.
Limitations
Our analysis only focuses on a single language model family (Pythia), in a a single language (English). To test the cross-lingual properties of adjective order, a wider range of language models should be examined. Our double adjective extraction method (§3.2) depends on the quality of the dependency parser that was used, and as such contains inaccuracies. Furthermore, a downside of using such a broad-coverage approach is that it becomes harder to draw targeted conclusions about the linguistic properties of cases on which the model fails or succeeds. Annotating cases containing properties that have been shown to be important for adjective order in the linguistics literature would provide deeper insights into model behavior on a linguistic level.
Appendix A Complementary Results
[…]
We provide a sample of CAP examples in Table 2, categorised by the contextual and isolated AOP scores of Pythia-12b.
[…]
CAP order always preferred
1. In solution, lead(II) hydroxide is a somewhat weak base, forming lead(II) ion, Pb, under weakly acidic conditions. $16.5$ $18.8$
2. It has the only deep water port in Cambodia. $16.9$ $15.4$
3. The greatest social reforms in Denmark are certainly the work of the last half-century. $15.4$ $16.3$
4. Icterids are unusual in songbirds because they have considerable sexual dimorphism. $16.9$ $14.5$
Swapped order always preferred
5. Couches are usually bought in a set together with cushions, which give them a bouncey and decorative comfortable touch. $-7.1$ $-6.8$
6. No, its 10 feet tall, and it’s a red big monster, demon like. $-7.3$ $-5.9$
7. It does this by following the grammatical basic rules of syntax. $-5.8$ $-6.0$
8. The daily average attendance from January through November 2010 was 22,133 people a day. $-5.8$ $-5.7$
Context improves AOP
9. Being a spherical 3-manifold, it is the only homology 3-sphere, besides the 3-sphere itself, with a finite fundamental group. $0.9$ $16.8$
10. Since the 2015 Styria municipal structural reform, it is part of the municipality Birkfeld. $-0.5$ $15.3$
11. I won it at a fair last night. $0.2$ $15.4$
12. I say that the man who does not play whist lays up a sad old age for himself. $2.6$ $17.5$
Context worsens AOP
13. This is an ability which may have been present in their last common ancestor in the Archaean. $12.3$ $4.0$
14. Jacob, you’ve had a complex partial seizure, which can cause psychosis, including religious psychosis. $7.3$ $0.3$
15. He was also a right-arm medium fast bowler with three wickets in test matches to his credit. $6.2$ $-0.6$
16. He wheeled round sharply, and distinguished her lying with helpless outspread arms on the couch. $4.0$ $-2.8$
Table 2: Most salient examples of the four quadrants in Figure 6.
A.2 Correlations to Cognitive Predictors
Correlations during training of AOP-$\Delta$ scores (Pythia-1.4b), with scores derived from various cognitive predictors.
A.3 Relative Number of Bigram Occurrences
Relative sizes of the 4 splits of CAP based on the number of bigram occurrences seen at each point in training. Note that an additional constraint on each split is that the swapped adjective order has not been seen at all, which explains why the 4 splits do not sum up to 1.
A.4 One-shot AOP
In Figure 7 we plot the AOP-$\Delta$ scores for the 27 adjective pairs that are seen exactly once in the first 10% of training.
A.5 Context Improvement Ratios
The relative number of items that are affected by context with respect to isolated AOP-$\Delta$ scores for each model size. These ratios correspond to the four quadrants of Figure 6A.
Figure 7: The AOP-$\Delta$ scores (Pythia-1.4b) across time for the subset of adjective pairs that are seen exactly once up to the 14,000th batch. Each curve is centered around the point where the adjective pair was encountered. The dotted line denotes the average AOP-$\Delta$ across all CAP items.

### v2

@@
-Black Big Boxes: Do Language Models Hide a Theory of Adjective Order?
+Black Big Boxes: Tracing Adjective Order Preferences in Large Language Models
 Abstract
-In English and other languages, multiple adjectives in a complex noun phrase show intricate ordering patterns that have been a target of much linguistic theory.
-These patterns offer an opportunity to assess the ability of language models (LMs) to learn subtle rules of language involving factors that cross the traditional divisions of syntax, semantics, and pragmatics.
-We review existing hypotheses designed to explain Adjective Order Preferences (AOPs) in humans and develop a setup to study AOPs in LMs: we present a reusable corpus of adjective pairs and define AOP measures for LMs.
-With these tools, we study a series of LMs across intermediate checkpoints during training.
-We find that all models’ predictions are much closer to human AOPs than predictions generated by factors identified in theoretical linguistics.
-At the same time, we demonstrate that the observed AOPs in LMs are strongly correlated with the frequency of the adjective pairs in the training data and report limited generalization to unseen combinations.
-This highlights the difficulty in establishing the link between LM performance and linguistic theory.
-We therefore conclude with a road map for future studies our results set the stage for, and a discussion of key questions about the nature of knowledge in LMs and their ability to generalize beyond the training sets.
-[^1] All code and data, including CAP, can be found here: github.com/jumelet/lm-adjorder
+In English and other languages, multiple adjectives in noun phrases follow intricate ordering patterns.
+These patterns have been widely studied in linguistics and provide a useful test case for assessing how language models (LMs) acquire graded and context-sensitive word order preferences.
+We ask to what extent adjective order preferences in LMs can be explained by distributional learning alone, and where models exhibit behaviour that goes beyond surface co-occurrence patterns.
+We find that LM predictions are largely explained by training data frequencies: simple $n$-gram statistics account for much of their behaviour and closely mirror the preferences learned during training.
+However, by analysing learning dynamics we reveal that models also generalize robustly to unseen adjective combinations, indicating that their behaviour cannot be reduced to memorization of observed orders alone.
+Moreover, we show how LMs leverage word order cues from sentence context, demonstrating with feature attribution methods that contextual cues are an additional driver of adjective order in LM output.
 1 Introduction
 Figure 1: We connect the adjective order preferences (AOP-$\Delta$, §3.1) of language models (here Pythia-12b) to the adjective order frequencies of the corpus they have been trained on (The Pile).
-We highlight various regions of interest: adjective pairs for which both orders are rare and that require the LM to generalize from other adjective orders; pairs for which one particular order is far more common that can be resolved from frequency alone; and orders with high variance that rely more strongly on context.
-Linguistic capabilities of language models (LMs) have been subject to a lot of research and benchmarking (e.g., Marvin and Linzen 2018; Hu et al.
-2020; Gauthier et al.
-2020).
-While LMs show impressive fluency in language generation and often score high on linguistic benchmarks, the mechanisms underlying their linguistic proficiency are hard to uncover.
-To what extent do LMs generalize beyond mere memorisation of training data?
-Do they acquire abstract linguistic rules and constraints that don’t boil down to basic co-occurrence patterns?
-How similar are linguistic generalizations in LMs to those that humans have?
-These questions have been asked and partially answered for different constructions and grammatical phenomena.
+We highlight various regions of interest: adjective pairs for which both orders are rare and that require the LM to generalize from other adjective orders; pairs for which one particular order is far more common that can be resolved from frequency alone; and orders with high variance.
+Language models show impressive fluency in language generation and often score high on linguistic benchmarks, but the mechanisms underlying their linguistic proficiency remain hard to uncover.
+In particular, it remains unclear to what extent a model’s knowledge of word order reflects memorization of surface co-occurrence patterns, and to what extent it supports more generalizable behaviour.
+What kinds of distributional and contextual information allow language models to go beyond observed word sequences when forming word order preferences?
+These questions have been posed and, in some cases, partially answered for various constructions and grammatical phenomena.
 The focus of such studies has mostly been on strict and binary grammatical constraints, where a violation results in an ungrammatical sequence.
 There has been less attention to how ‘softer’ linguistic constraints are captured by LMs.
-Such constraints can give rise to slight and defeasible preferences rather than sharp acceptability contrasts.
-One example is adjective order in noun phrases where multiple adjectives modify one noun, e.g.
-large wooden box.
-Changing the order of adjectives does not necessarily lead to unacceptability, but can result in a slight decrease in naturalness (Dyer et al., 2023) (‘>’ marks preference):
-  1. large wooden box $>$ wooden large box
-Moreover, these preferences are context-dependent: often, a context can be found where preferences switch – for instance, when a contrasting property is introduced (Teodorescu, 2006).
+Such constraints can give rise to slight and defeasible preferences.
+One example is adjective order in noun phrases where multiple adjectives modify one noun.
+Changing the order of adjectives does not necessarily lead to unacceptability, but can result in a slight decrease in naturalness.
+E.g., (‘>’ marks preference):
+large wooden box >> wooden large box
+Moreover, these preferences are context-dependent: often, a context can be found where preferences switch – for instance, when a contrasting property is introduced (Teodorescu, 2006):
   1. Take a wooden large box, not a plastic one.
-  2. $>$   … large wooden box …
-These adjective order preferences (AOPs) have been subject to a lot of interest in theoretical linguistics, but proved difficult to capture due to their graded and context-sensitive nature (see Levshina et al.
-2023 for arguments in favor of a systematically gradient approach to word order).
-A number of factors ranging from superficial properties like adjective length to semantic and pragmatic properties to information-theoretic properties of a sentence have been identified as predictors for AOP.
-But even with all those factors combined together, a large portion of AOPs remains unaccounted for: the best model combining different predictors explains only 71% of human AOPs (Dyer et al., 2023).
-Adjective order offers an opportunity to assess the ability of LMs to learn subtle linguistic preferences involving implicit, hard-to-identify factors.
-This paper is the first study of AOP in LMs.
-We put together metrics and data, and conduct a series of experiments that explore AOP in LMs, its relation to properties of training data and the extent to which LMs learn more abstract principles behind AOP, generalizing to unseen items.
+  2. >> … large wooden box …
+Such adjective order preferences (AOPs) have been subject to extensive study in theoretical linguistics, but have proved difficult to capture due to their graded and context-sensitive nature.
+A wide range of factors have been proposed as predictors of AOPs, ranging from superficial properties like adjective length to semantic and pragmatic properties to information-theoretic properties.
+Even when all those factors are combined, however, a significant portion of AOPs remains unaccounted for: the best-performing multi-factor models explain only about 71% of human AOPs (Dyer et al., 2023).
+This gap highlights the complexity of adjective order preferences and motivates their use as a challenging test case for studying how language models form word order preferences.
+Rather than treating language models as explanatory accounts of adjective order, we analyse their behaviour in order to ask which aspects of their performance can be traced to distributional properties of the training data, and where additional effects such as generalization beyond observed data and sensitivity to sentence context arise.
+We conduct a series of experiments that explore AOP in LMs, its relation to properties of training data statistics, the extent to which LMs generalize beyond memorization, and how sentence context shapes model preferences.
 The contributions of our paper are the following:
-[…]
-  • We show that LMs are impressively good predictors of naturally occurring adjective orders, with AOP prediction accuracy as high as 94.1%.
-We identify three stages of AOP acquisition in LMs that lead to their performance at the end of the training (§4).
-  • We demonstrate that linguistic and cognitive predictors only partly account for LM performance in AOP prediction.
-We identify the gap between linguistic theory and LM performance in the AOP domain that needs to be explained (§5).
-  • To pinpoint the role of training data co-occurrence statistics in LM AOP predictions, we study the Pile (Gao et al., 2020) – a large-scale corpus that Pythia models were trained on.
-Simple bigram statistics turn out to be a powerful AOP predictor for naturally occurring adjective orders on their own, at 90.3% accuracy (§6).
-Figure 1 summarizes a close relation between count statistics and LM predictions, questioning whether LM predictions are significantly driven by anything beyond $n$-gram counts.
-  • At the same time, we find two ways in which LMs go beyond simple collocational statistics: 1) We test LM predictions for adjectival combinations not seen during training (a challenging setup due to the large size of the Pile; we overcome this obstacle by turning to intermediate checkpoints and corresponding portions of training data).
-We find some—but limited—generalization to new data (§6); 2) We study the effect of context on AOP.
-We conclude that the sentence helps LMs make better AOP predictions compared to noun phrase with adjectives taken in isolation or in a set of random contexts, which indicates that AOP mechanisms in LMs go beyond $n$-gram counts in training data (§7).
+  • We provide the first systematic empirical analysis of adjective order preferences (AOPs) in large language models, showing that models trained on natural language data exhibit strong and consistent preferences for naturally occurring adjective orders.
+Using intermediate training checkpoints from the Pythia model suite, we characterize how these preferences emerge over the course of training and identify three distinct stages of acquisition (§4).
+  • To assess the role of training data statistics, we analyse $n$-gram distributions from the Pythia training corpus using Infini-gram (Liu et al., 2024).
+We show that simple bigram statistics account for a large portion of AOP behaviour, and are strongly correlated with model predictions (§5.1 and Figure 1).
+  • We then assess the extent to which model behaviour generalizes beyond memorization by evaluating AOP predictions for adjective combinations not seen during training.
+Despite the strong influence of distributional statistics, models exhibit robust generalization to unseen adjective orders (§5.2).
+  • Finally, we investigate the role of sentence context in shaping adjective order preferences.
+We show that contextual information systematically improves AOP predictions relative to neutral or random contexts, and use feature attribution methods to identify two sources of contextual influence: local collocational cues and earlier semantic cues that bias adjective ordering (§6).
+Altogether, our experiments provide a behavioural characterization of how adjective order preferences in language models relate to training data statistics, generalization beyond memorization, and contextual information, and offer a general framework for studying word order in language models.
 2 Background
@@
 Furthermore, corpus studies have shown that many counterexamples against various versions of these hierarchies can be found (Truswell, 2009), demonstrating the need for a more fine-grained analysis (Trotzke and Wittenberg, 2019).
-One line of work looks for semantic driving forces of AOP, placing emphasis on adjectival properties like intersectivity and subsectivity (Truswell, 2009) or identifying conditions under which adjective order constraints can be lifted, such as focus (Vendler, 1968; Teodorescu, 2006).
+One line of work looks for semantic driving forces of AOPs, placing emphasis on adjectival properties like intersectivity and subsectivity (Truswell, 2009) or identifying conditions under which adjective order constraints can be lifted, such as focus (Vendler, 1968; Teodorescu, 2006).
 Other approaches take a broader perspective, leveraging insights from quantitative linguistics, distributional semantics and information theory, and connecting this to psycholinguistic theories of human language processing.
@@
 This follows from a broader theory of language production, according to which more accessible and familiar material is produced first (Bock, 1982; Ferreira and Dell, 2000) and is comprehended faster (Arnon and Snider, 2010).
-Collocational patterns of adjectives and nouns have been shown to be highly predictive of adjective order: adjectives most closely connected to the noun in term of co-occurrence appear closer to it (Sweet, 1898; Byrne, 1979; Lapata et al., 1999).
+Collocational patterns of adjectives and nouns have been shown to be highly predictive of adjective order: adjectives most related to the noun appear closer to it (Sweet, 1898; Byrne, 1979; Lapata et al., 1999).
 This intuition has been operationalized through pointwise mutual information (PMI) (Hahn et al., 2018; Futrell et al., 2020a), which aligns with more general theories of word order arguing that distance between words that closely predict each other tends to be small (Gibson, 1998; Futrell et al., 2020b).
@@
 Although the majority of these approaches focus on English, various cross-lingual accounts have demonstrated their universal nature: adjective order tends to be driven by the same principles across languages (Wulff and Gries, 2015; Leung et al., 2020; Scontras, 2023; Dyer et al., 2023).
-Finally, various papers highlight the need for a multi-factorial account of adjective order: it is unlikely adjective order is driven by a single factor and therefore it’s the result of various competing pressures (Wulff, 2003; Trotzke and Wittenberg, 2019; Dyer et al., 2023).
+Finally, various papers highlight the need for a multi-factorial account of adjective order: it is unlikely that adjective order is driven by a single factor, and therefore it results from competing pressures (Wulff, 2003; Trotzke and Wittenberg, 2019; Dyer et al., 2023).
 Relatedly, a recent position paper (Levshina et al., 2023) argues for a fundamentally gradient theory of word order that would be able to capture its hybrid and multi-faceted nature.
-In our work, we compare LM predictions to various adjective order theories.
-Additionally, we examine the utility of LMs for investigating the role that context plays on word order: since almost all theories focus on adjectives in isolation, LMs may provide a fruitful starting point for shaping a contextual account of adjective order.
+In our work, we study adjective order preferences in large language models as a means of analysing how such models handle graded and context-sensitive linguistic phenomena.
+Rather than evaluating language models as alternatives to existing linguistic theories of adjective order, we use them as a tool for probing which sources of information influence adjective ordering decisions.
+In particular, while most theoretical accounts of adjective order focus on adjectives in isolation, language models naturally condition on rich sentential context.
+This makes them a useful testbed for investigating how contextual cues shape adjective order preferences.
+By analysing the contextual information that influences model predictions, we aim to identify patterns that may inform future empirical studies of adjective order and the role of context in human language use.
 2.2 Word Order in Language Models
@@
 The importance of word order in LMs has been a topic of debate, with various works claiming that downstream performance is not affected by scrambled inputs (Malkin et al., 2021; Sinha et al., 2021), although it has been shown that LMs are able to retain a notion of word order through their positional embeddings (Abdou et al., 2022).
-It has been argued that LMs acquire an abstract notion of word order that goes beyond mere $n$-gram co-occurrence statistics (Futrell and Levy, 2019; Kuribayashi et al., 2020; Merrill et al., 2024), a claim that we in this paper assess for large-scale LMs in the context of adjective order.
-Finally, numerous works have investigated the trade-off between memorization and generalization in LMs: it has been shown that larger LMs are able to memorize entire passages from the training data (Biderman et al., 2023a; Lesci et al., 2024; Prashanth et al., 2024), but generalization patterns for grammatical phenomena have also been shown to follow human-like generalization (Dankers et al., 2021; Hupkes et al., 2023; Alhama et al., 2023).
+It has been argued that LMs acquire a notion of word order that generalizes beyond $n$-gram co-occurrence statistics (Futrell and Levy, 2019; Merrill et al., 2024; Michaelov et al., 2025; Wang et al., 2025), a claim that we in this paper assess for large-scale LMs in the context of adjective order.
+Finally, numerous works have investigated the trade-off between memorization and generalization in LMs: it has been shown that larger LMs are able to memorize entire passages from the training data (Biderman et al., 2023a; Lesci et al., 2024; Prashanth et al., 2024; van der Wal et al., 2025), but generalization patterns for grammatical phenomena have also been shown to follow human-like generalization (Dankers et al., 2021; Hupkes et al., 2023; Alhama et al., 2023).
 3 Methods
@@
 For a double adjective phrase $\textsc{a}_{1}\textsc{a}_{2}\textsc{n}$ that is extracted from a corpus with context c, we compute a model’s preferred order as follows:
-[…]
-[…]
+$\textup{AOP-}\Delta(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})$ $=P(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})-P(\textsc{a}_{2}\textsc{a}_{1}\textsc{n})$ (1)
+$\textup{AOP-}\Delta(\textsc{a}_{1}\textsc{a}_{2}\textsc{n}|\textsc{c})$ $=P(\textsc{a}_{1}\textsc{a}_{2}\textsc{n}|\textsc{c})-P(\textsc{a}_{2}\textsc{a}_{1}\textsc{n}|\textsc{c})$ (2)
 All probabilities are in log space.
@@
 Misra and Mahowald (2024) consider a similar metric for measuring properties of word order preferences in LMs, based on corrupted word orders.
-[…]
-[…]
+3.2 Evaluation Corpus
+To evaluate adjective order preferences across a wide range of adjective pairs and contexts, we make use of the double adjective corpus of Jumelet et al.
+(2025).
+In our experiments we refer to this corpus as CAP: Corpus of Adjective Pairs.
+Earlier work has extracted adjective pairs from the Universal Dependencies treebank directly (Dyer et al., 2023), but we instead use CAP to collect model judgments across a wider range of adjective pairs.
 Procedure
-As a starting point, we use the BabyLM 10M corpus (Warstadt et al., 2023), which consists of a mix of various English language sources including Wikipedia articles, books, and transcribed dialogue data.
-From this corpus we make a pre-selection of sentences containing multiple adjectives by filtering all sentences containing at least two adjectives, based on a list of 11,000 adjectives released by Futrell et al.
+The evaluation corpus is derived from the BabyLM 100M corpus (Charpentier et al., 2025), which consists of a mix of various English language sources including Wikipedia articles, books, and transcribed dialogue data.
+From this corpus a pre-selection of sentences containing multiple adjectives is made by filtering all sentences containing at least two adjectives, based on a list of 11,000 adjectives released by Futrell et al.
 (2020a).
-Next, we generate dependency parse trees for each sentence using spaCy and select all sentences in which a noun contains two amod relations to an adj that is part of the adjective list.
-This procedure results in 2678 double adjective sentences.
-We provide a sample of CAP sentences in Appendix A.1.
-From these sentences we also generated corrupted sentences in which the adjective order is swapped.
-In case the first adjective is preceded by an indefinite article we ensure the article is changed accordingly to satisfy morpho-phonological constraints.
-Here is an example of an extracted phrase and its swapped counterpart:
-  1. $+$
+Next, dependency parse trees are generated for each sentence using spaCy’s RoBERTa-based parser.
+All sentences are then selected in which a noun contains two amod relations to an adj that is part of the adjective list.
+This procedure results in 9396 double adjective sentences, consisting of 2341 unique adjectives.
+A sample of sentences from the corpus is provided in Appendix A.1.
+From these sentences we generate corrupted sentences in which the adjective order is swapped.
+In case the first adjective is preceded by an indefinite article we ensure the article is changed accordingly to satisfy morpho-phonological constraints, for example:
+  1. ++
     Peter received a full athletic scholarship [..]
-  2. --
+  2. -
     Peter received an athletic full scholarship [..]
+We treat the corpus-attested order as a proxy for canonical preference, while acknowledging that adjective order preferences are graded and context-dependent
+Figure 2: A–B: AOP-% and AOP-$\Delta$ scores for Pythia models of increasing size.
+C–D: AOP-% and AOP-$\Delta$ scores for Pythia-1.4b during training.
+We highlight the three learning phases: 1) initialization, 2) acquisition, and 3) consolidation.
 3.3 Models
@@
 Importantly, all intermediate checkpoints are available for these models in logarithmic intervals, which allows us to investigate the learning dynamics at a fine-grained level.
-Figure 2: A–B: AOP-and AOP-$\Delta$ scores for Pythia models of increasing size.
-C–D: AOP-and AOP-$\Delta$ scores for Pythia-1.4b during training.
-We highlight the three learning phases: 1) initialization, 2) acquisition, and 3) consolidation.
-4 Experiment 1: AOP in LLMs
-In this first set of experiments we evaluate a series of pretrained LLMs on CAP using the AOP-$\Delta$ and AOP-metrics of §3.1.
+4 AOP in LMs
+In this first set of experiments we evaluate a series of pretrained LMs on CAP using the AOP-$\Delta$ and AOP-% metrics of §3.1.
 Model Size
-For the eight Pythia models of increasing size we plot the AOP-performance in Figure 2A, and the AOP-$\Delta$ performance in Figure 2B.
-It can be seen that performance improves with model size: contextual AOP-performance increases from 89.6% for the 70m model to 94.1% for the 12b model.
-AOP-does not increase consistently with size, however, from the 1b model onward it reaches a plateau.
+For the eight Pythia models of increasing size we plot the AOP-% performance in Figure 2A, and the AOP-$\Delta$ performance in Figure 2B.
+It can be seen that performance improves with model size: contextual AOP-% performance increases from 85.0% for the 70m model to 91.7% for the 12b model.
+AOP-% does not increase consistently with size, however, from the 1b model onward it reaches a plateau.
 Interestingly, AOP-$\Delta$ does increase consistently with size: this shows that larger models become more certain about their AOP judgment.
@@
 From these plots we can identify three distinct learning phases: 1) an initialization phase in which AOP judgments have not been formed yet, 2) an acquisition phase in which AOP judgments are rapidly being formed, and 3) a consolidation phase in which AOP judgments remain stable and only the role of context is being reinforced.
-A similar three-phase learning dynamic has been observed by Van Der Wal et al.
+A similar tri-phase learning dynamic has been observed by Van Der Wal et al.
 (2022) and Chen et al.
@@
 Note that the consolidation phase is reached relatively quickly: after around 2000 batches, which is ~1.4% of the total model training.
-This demonstrates that adjective order is a relatively low-level linguistic feature that is acquired early during training.
+This suggests that sensitivity to adjective order preferences emerges early during model training.
 Furthermore, it highlights the importance of investigating learning dynamics at a logarithmic scale: with linear checkpoints we would not obtain this fine-grained level of precision.
@@
 In our experiments so far we have measured AOP based on the probabilities of the full adjective-noun triplet.
-Preference order can arise at different locations of the phrase.
-The most natural location for this is the probability of the second adjective after having seen the first adjective, but context could also play a role for preferring which adjective comes at the first position, and a noun could be more probable after having seen two adjectives in their natural order.
-In Figure 3 we plot the average token probabilities on CAP for Pythia-12b, in isolation and contextual.
+Order preference can arise at different locations of the phrase.
+The most natural location for this is the second adjective after having seen the first adjective, but context could also play a role for preferring which adjective comes at the first position, and a noun could be more probable after having seen two adjectives in their natural order.
+In Figure 3 we plot the average token probabilities on CAP for Pythia-12b, in isolation and in context.
 We observe that, as expected, the largest probability difference occurs at the position of the second adjective: the model is highly surprised by encountering the second swapped adjective after having seen the first.
@@
 The Pythia LMs acquire a strong sense of adjective order that is learned early in training at around 1.5% of the total amount of training.
-We showed that sentence context has a positive effect on AOP, bumping up AOP-from 91% to around 94%.
+We showed that sentence context has a positive effect on AOP, bumping up AOP-% from 88% to around 91%.
 By localizing the AOP scores on the token-level we show how these order preferences manifest at different points in a phrase.
-In the next experiments we will investigate these results in more detail, connecting LM AOP to cognitive predictors and training data statistics.
-5 Experiment 2: Cognitive Predictors
+In the next experiments we will investigate these results in more detail, connecting LM AOPs to training data statistics.
+5 The Role of Training Data Statistics in AOP
+Having established that language models achieve high accuracy on adjective order prediction, we now examine the extent to which this behaviour can be explained by distributional properties of the training data.
+Specifically, we ask to what extent simple co-occurrence statistics explain model predictions, and where systematic deviations from such statistics arise.
+To ensure that our subsequent analyses are not specific to a single model family, we first examine adjective order behaviour across a range of contemporary large language models.
+Table  5 shows the performance of four contemporary, mid-size, open-source LLMs: Llama3-8b, OLMo2-7b, OLMo3-7b and Qwen3-8b (Grattafiori et al., 2024; Yang et al., 2025; OLMo et al., 2025a, b).
+These models perform similarly to Pythia (slightly worse, at 89 to 91%) with high AOP-$\Delta$ correlation, supporting the use of the Pythia models as a representative case in our subsequent analyses.
+As a next step, we assess whether LM adjective order behaviour can be accounted for by cognitive predictors proposed in prior work (§2.1).
+Compared to such predictors, language models exhibit adjective order behaviour that is not well explained by any single factor.
+In Table 5 we show the accuracy of two predictors: PMI between adjective and noun (62.4%), and adjective subjectivity (69.3%).
+As show in Appendix A.2, these predictors show limited overlap with the AOP-$\Delta$ scores, indicating that LM behaviour cannot be straightforwardly reduced to the effects captured by these individual predictors.
 […]
-Metric AOP-$\rho(\textup{AOP-}\Delta(\bullet))$ $\rho(\textup{AOP-}\Delta(\bullet|c))$
-AOP-$\Delta(\bullet)$ 91.6 0.71
-[…]
-[…]
-PMI(a;n) 62.4 0.27 0.20
-Subj(a) 69.3 0.07 0.08
-[…]
-%$\%(\textsc{a}_{1}\textsc{a}_{2})$ 90.3 0.58 0.53
-%$\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})$ 87.6 0.54 0.53
-Table 1: Predictive accuracy of adjective order on CAP for various metrics: AOP-$\Delta$ for Pythia-12b, 3 cognitive predictors (§5), and relative n-gram counts from The Pile (§6).
+Table 1: Predictive accuracy of adjective order on CAP for various metrics: AOP-$\Delta$ for Pythia-12b and relative $n$-gram counts from The Pile (§5).
 The highest accuracy for each group is bolded.
 For each metric we provide the Spearman correlations with respect to the AOP-$\Delta$ scores of Pythia-12b.
-As discussed in §2.1, a range of cognitively plausible predictors have been proposed to explain adjective order.
-In the previous section we have shown that LMs are highly capable of predicting adjective order, and in this section we investigate how their predictions are related to various theories on adjective order.
-We evaluate the performance of three cognitive predictors—Length, PMI, and Subjectivity— and correlate them against AOP-$\Delta$ scores across training.
-5.1 Cognitive Predictors
-For each predictor we compute its score for a particular adjective pair. An adjective order predictor $f$ returns a graded score for adjective order, akin to our formulation of AOP-$\Delta$.
-Length
-For Length—which hypothesizes that the first adjective is the shortest—the score for adjectives a1 and a2 is expressed as
-$f(\textsc{a}_{1},\textsc{a}_{2})=|\textsc{a}_{2}|-|\textsc{a}_{1}|$
-We also considered expressing length as the number of tokenizer subwords, since this is the only notion of length LMs have access to, but found that this led to similar results as string length.
-PMI
-We follow the procedure of Futrell et al. (2020a) for computing PMI, which is based on the distribution over adjective relations in dependency parse trees. It is defined as
-$\textup{PMI}(\textsc{a};\textsc{n})=P(\textsc{a},\textsc{n})/(P(\textsc{a})\cdot P(\textsc{n}))$
-$P(\textsc{a},\textsc{n})$ is computed as the normalized number of occurrences of adjective a and noun n in amod relation.
-$P(\textsc{a})$ and $P(\textsc{n})$ are obtained by marginalizing over this joint distribution.
-PMI hypothesizes that adjectives closer to the noun have a higher PMI:
-$f(\textsc{a}_{1},\textsc{a}_{2},\textsc{n})=\textup{PMI}(\textsc{a}_{2};\textsc{n})-\textup{PMI}(\textsc{a}_{1};\textsc{n})$
-Subjectivity
-We leverage the 450 subjectivity ratings provided by Dyer et al.
-(2023), to compute subjectivity scores for the subset of adjective pairs that contain a rating.
-This procedure retains 37.5% of the adjective pairs in CAP (1005 items).
-Subjectivity as predictor hypothesizes that more subjective adjectives occur first (Scontras et al., 2017):
-$f(\textsc{a}_{1},\textsc{a}_{2})=\textup{Subj}(\textsc{a}_{1})-\textup{Subj}(\textsc{a}_{2})$
-5.2 Predictive Accuracy
-We present the predictive accuracy of each predictor in Table 1.
-Subjectivity is the strongest predictor with 69.3% accuracy, followed by PMI (62.4%) and Length (49.6%).
-While Subjectivity reaches the highest accuracy, its correlation with the LM’s AOP-$\Delta$ scores is surprisingly the lowest of the three.
-We plot these correlations over time in Appendix A.2: all three correlations develop in the acquisition phase and remain stable in the consolidation phase.
-We also compute the overlap between these three predictors on the 1005 items for which we have subjectivity scores:
-[Uncaptioned image][Uncaptioned image]
-The first Venn diagram makes clear that the cognitive predictors have low agreement: only 213 orders are predicted correctly by all three.
-The broad coverage (93.1%) of the union, however, shows that most of the CAP items are explained by at least one of these predictors.
-In the four diagrams on the right we show the overlap of the predictors with the AOP-$\Delta$ scores of Pythia-12b.
-All three predictors and their intersection are almost fully subsumed by the AOP-$\Delta$ predictions.
-This demonstrates that AOP-$\Delta$’s unique contribution goes significantly above that of the ensemble of the cognitive factors, raising the question whether LMs ‘hide’ a meaningful linguistic theory that extends beyond established factors.
-6 Experiment 3: Is AOP Driven by Data Statistics?
-Figure 4: Correlations during training of LM probabilities for single adjectives, adjective pairs, and adjective-noun triplets with respect to their frequency in The Pile.
-In order to truly understand what LMs base their adjective order preferences on, we need to get a better insight into how double adjective constructions are distributed in the training data.
-The way that contemporary LMs are trained nowadays makes this challenging: many models are released without open access to their training data, and the scale of training corpora that are available poses a challenge for targeted analysis.
-6.1 WIMBD
-To overcome this, we make use of the WIMBD API (What’s In My Big Data, Elazar et al., 2023), which provides a fully indexed search engine over various large-scale corpora such as The Pile, on which the Pythia models have been trained.
-WIMBD makes it possible to retrieve the exact $n$-gram statistics that LMs have been trained on, and as such provides a detailed insight into the collocational factors that play a direct role in shaping AOP.
+AOP scores for the four LLMs denote contextual AOP.
+5.1 Infini-gram
+In order to better understand the factors underlying LM adjective order behaviour, we examine how double adjective constructions are distributed in the training data.
+The way that LMs are trained nowadays makes this challenging: many models are released without open access to their training data, and the scale of training corpora that are available poses a challenge for targeted analysis.
+To overcome this, we make use of the Infini-gram API (Liu et al., 2024), which provides a fully indexed search engine over various large-scale corpora such as The Pile, on which the Pythia models have been trained.
+Infini-gram makes it possible to retrieve the exact $n$-gram statistics that LMs have been trained on, and as such provides a detailed insight into the collocational factors that play a direct role in shaping AOP.
 We retrieve the uni-, bi- and trigram counts for all the adjective-noun triplets in CAP for The Pile, as well as for the swapped orders.
@@
 We express these relative $n$-gram counts as
-%$\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})=\log\#(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})-\log\#(\textsc{a}_{2}\textsc{a}_{1}\textsc{n})$
-with similar formulations for adjective bigrams (%($\textsc{a}_{1}\textsc{a}_{2}$)) and unigrams (%($\textsc{a}_{1}$)).
+$\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})=\log\#(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})-\log\#(\textsc{a}_{2}\textsc{a}_{1}\textsc{n})$
+[…]
 Predictive Accuracy
 We present the predictive accuracy of these relative $n$-gram occurrences in Table 5.
-%(a1) predicts adjective order with 60.5%, which provides weak evidence that more frequent adjectives tend to occur first (Martin, 1969; Trotzke and Wittenberg, 2019).
-%(a1a2) is the strongest predictor of adjective order—except for the LMs—at 90.3%.
+%(a1) predicts adjective order with 55.2%, which provides weak evidence that more frequent adjectives tend to occur first (Martin, 1969; Trotzke and Wittenberg, 2019).
+%(a1a2) is the strongest predictor of adjective order at 86.8%.
 This shows that a large majority of the items in CAP can be resolved based on bigram statistics alone, which is leveraged by the LMs during training.
-However, n-gram collocations alone do not fully explain LM performance: a large amount of trigrams are not part of the Pile and performance for %$\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})$ drops to 87.6%.
+However, $n$-gram collocations alone do not fully explain LM performance: a large amount of trigrams are not part of the Pile and performance for $\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})$ drops to 76.7%.
+Figure 4: Correlations during training of LM probabilities for single adjectives, adjective pairs, and adjective-noun triplets with respect to their frequency in The Pile.
 Correlations
-To investigate how n-gram statistics play a role in the formation of AOP-$\Delta$ during training, we compute the Spearman correlation between various log-transformed n-gram counts and a LM’s AOP-$\Delta$ on intermediate checkpoints.
+To investigate how $n$-gram statistics relate to the formation of AOP-$\Delta$ during training, we compute the Spearman correlation between various log-transformed $n$-gram counts and a LM’s AOP-$\Delta$ on intermediate checkpoints.
 We plot this in Figure 4, for contextual and isolated AOP-$\Delta$.
@@
 Then, at the start of the acquisition phase (2), in which AOP performance rapidly improves, the model starts to incorporate linguistic context in its uni- and bigram predictions.
-Finally, in the consolidation phase (3), trigram predictions start being affected by context as well, whereas unigram and bigram predictions remain relatively stable.
+Finally, in the consolidation phase (3), trigram predictions start being affected by context as well, whereas unigram and bigram predictions remain stable.
 The fully trained model strongly reflects these correlations to data statistics, which explains the strong relation we observed in Figure 1 between bigram frequencies and AOP-$\Delta$.
+This successive alignment from unigram to higher $n$-gram statistics has also been observed in prior work (Chang and Bergen, 2022; Michaelov et al., 2025).
 5.2 Generalizing to Unseen Adjective Orders
-Figure 5: The contextual AOP-performance for Pythia-70m and 1.4b across training, split out for items that have been seen 0, 1, 2 to 10, and more than 10 times at each specific checkpoint.
-Having access to the n-gram distribution a model was trained on allows us to investigate its performance on unseen adjective orders: is the model able to generalize from an abstract signal of adjective order to novel orders?
-However, due to the enormous size of The Pile (300B tokens) it turns out that number of unseen adjective pairs in CAP is too small to draw significant conclusions from (only 4 out 2678 pairs).
+Having access to the $n$-gram distribution a model was trained on allows us to investigate its behaviour on adjective orders not previously observed during training.
+We ask whether models exhibit systematic generalization to unseen adjective combinations, beyond direct memorization of observed co-occurrence patterns.
+However, due to the enormous size of The Pile (300B tokens) it turns out that number of unseen adjective pairs in CAP is too small to draw significant conclusions from (only 39 out 9396 pairs, 0.4%).
 This leaves us with two options: either we collect an additional sample of rare adjective pairs that are not part of the Pile, or we investigate the model at intermediate checkpoints where a greater amount of adjective pairs has not been encountered yet.
@@
 We therefore go for the second option, and set up a procedure to measure the data statistics during training.
+Figure 5: The contextual AOP-% performance for Pythia-70m and 1.4b across training, split out for items that have been seen 0, 1, 2 to 10, and more than 10 times at each specific checkpoint.
+We provide the size of these splits across training in Appendix A.3.
 Intermediate $n$-gram Counts
-Since WIMBD only provides $n$-gram statistics for The Pile as a whole, we need to collect the intermediate statistics ourselves.
-Using the Pythia batch viewer (Biderman et al., 2023b) this becomes possible: our implementation is able to process the first 10% of The Pile (up to batch 14,000; 29.4B tokens) in around 3 hours.
-We collect the bi-gram counts of all adjective pairs in CAP on the batch level, allowing us to determine exactly when and how often a specific adjective pair has been seen by the model.
+Since Infini-gram only provides $n$-gram statistics for The Pile as a whole, we need to collect the intermediate statistics ourselves.
+Using the Pythia batch viewer (Biderman et al., 2023b) this becomes possible: our implementation is able to process the first 10% of The Pile (up to batch 14,000; 29.4B tokens) under one hour on a consumer-grade laptop.
+We collect the bi-gram counts of all adjective pairs in CAP on the batch level, allowing us to determine exactly when and how often a specific adjective pair has been observed during training.
 Unseen Adjective Pairs
-Based on these intermediate $n$-gram counts, we compute the AOP-scores for various splits based on the number of times an adjective pair has been encountered.
-We focus on the cases where the swapped order has never been encountered: this allows us to isolate the moment a model encounters an adjective combination for the first time, and measure its ‘zero-shot’ performance on cases that are not already skewed towards to swapped order.
+Based on these intermediate $n$-gram counts, we compute the AOP-% scores for various splits based on the number of times an adjective pair has been encountered.
+We focus on the cases where the swapped order has never been encountered: this allows us to isolate the moment a model encounters an adjective combination for the first time, and measure its ‘zero-shot’ performance on these cases.
 For each model checkpoint we collect the $n$-gram counts at that particular point in training, and split the CAP items based on the number of times an a1a2 bi-gram has been seen.
-We provide the size of these splits across training in Appendix A.3.
 We present the results for Pythia-70m and 1.4b in Figure 5.
-The performance on unseen pairs is remarkably high, reaching around 85% at the highest point for both models.
-Interestingly, the gap between unseen pairs and pairs seen once is far greater for the 1.4b model than the small 70m model.
-This shows that this larger model is able to acquire adjective order preference faster: a single exposition to an order immediately leads to an increase in AOP-$\Delta$.
+The performance on unseen pairs is remarkably high, reaching around 85% at the highest point.
 We provide a more detailed item-level analysis for these items in Appendix A.4: while it is clear on the split level that a single occurrence improves AOP-%, we do not observe such a clear effect on the item level.
 Interestingly, as the number of unseen adjective pairs decreases, the performance on unseen pairs decreases as well.
-This shows that the predicting the order for the set of unseen pairs around batch 2,000 (6% of CAP) could be better inferred from a general signal of adjective order than the set at batch 14,000 (1.3% of CAP).
+As training progresses and the pool of unseen adjective pairs becomes smaller, performance on these unseen items decreases.
+Since the set of unseen pairs at later checkpoints is a strict subset of earlier unseen sets, this trend likely reflects increasing item difficulty rather than a qualitative change in model generalization.
 Conclusions
-The question we set out to answer in this set of experiments is whether AOPs are driven by data statistics.
-We approached this question both in relation to the $n$-gram distribution of the Pile and to the shifting $n$-gram distribution across training.
-On the one hand, AOP is strongly driven by data statistics: we find a strong correlation between AOP-$\Delta$ and count-based AOP scores, which is also visualized in Figure 1.
-However, the relatively strong performance on the set of unseen adjective pairs during training indicates that LMs are able to infer order from a more general signal for adjective pairs it has not encountered before.
-7 Experiment 4: The Role of Context
+The question we set out to address in this set of experiments is the extent to which adjective order behaviour in language models can be explained by direct co-occurrence statistics.
+We approached this question by analysing the overall $n$-gram distribution of the Pile and its evolution over the course of training.
+On the one hand, AOP is strongly connected to training data statistics: we observe a strong correlation between AOP-$\Delta$ and count-based AOP scores, as visualized in Figure 1.
+On the other hand, the relatively strong performance on adjective pairs not previously observed during training indicates that models generalize adjective order behaviour beyond direct memorization of seen co-occurrences.
 Figure 6: A: Item-level AOP-$\Delta$ scores with and without sentence-context for Pythia-12b.
 We provide a sample of sentences from each quadrant in Appendix A.1.
-B: The relative impact ($c$-$\Delta$, Eq.
+B: The contextual gain ($c$-$\Delta$, Eq.
 5) of an item’s true context compared to an expectation over random contexts (Eq.
 4).
+6 The Role of Context
 In §4 we established that LMs have a thorough grasp of adjective order that manifests itself early during training.
 We showed that having access to sentence context boosts performance, and we now investigate the role of context in more detail.
-Contextual Improvements
-To gain a better insight how much items are impacted by context, we investigate AOP-$\Delta$ scores on an item-level with and without context.
+6.1 Quantifying Context Impact
+Context improves predictions a lot
+To understand in what ways AOP is impacted by context, we investigate AOP-$\Delta$ scores on an item-level with and without context.
 In Figure 6A, we plot the AOP-$\Delta$ scores in isolation against the contextual scores, and mark the four quadrants of items that result in a positive or negative score after adding context.
-In the majority of cases (89.4%), AOP-$\Delta$ is positive both with and without context: the model is able to determine adjective order based on the adjective-noun triplet alone.
-In 4.8% of the cases, however, adding context flips a negative AOP-$\Delta$ score to a positive one.
-Adding context has a boosting effect in general: in 65.1% of the cases, AOP-$\Delta$ is increased by adding context.
-We provide an overview of how these ratios are affected by model size in Appendix A.5: the number of cases where context improves AOP-is actually larger for small models, but so is the number of cases where context worsens it.
+In the majority of cases (85.2%), AOP-$\Delta$ is positive both with and without context: the model is able to determine adjective order based on the adjective-noun triplet alone.
+In 6.5% of the cases, however, adding context flips a negative AOP-$\Delta$ score to a positive one.
+Adding context has a boosting effect in general: in 67.6% of the cases, AOP-$\Delta$ is increased by adding context.
+We provide an overview of how these ratios are affected by model size in Appendix A.5: the number of cases where context improves AOP-% is actually larger for small models, but so is the number of cases where context worsens it.
+Figure 7: Contextual AOP-% and Context Gain against Sentence Length and (Adjective) Order Entropy for all Pythia models.
+Scores are binned based on quantiles: $x$-tick labels indicate the lower bound of each bin (e.g.
+$[3,6)$ for the first tick).
 Random Contexts
@@
 To test this, we compute an expectation over AOP-$\Delta$ scores with random contexts, in which an adjective pair is preceded by a randomly sampled context from another adjective pair in CAP:
+$\textup{AOP-}\Delta(\textsc{a}_{1}\textsc{a}_{2}\textsc{n}\mid\mathcal{C})=\mathbb{E}_{c\sim\mathcal{C}}\left[\textup{AOP-}\Delta(\textsc{a}_{1}\textsc{a}_{2}\textsc{n}\mid c)\right]$ (4)
+We compute the expectation over a random sample of 500 contexts, taken from CAP ($\mathcal{C}$).
+To express the relative impact of adding context, we regress out a model’s isolated AOP-$\Delta$ score by subtracting it from the contextual scores, which we call the contextual gain ($c\textup{-}\Delta$):
+$c\textup{-}\Delta(\bullet|c)=\textup{AOP-}\Delta(\bullet|c)-\textup{AOP-}\Delta(\bullet)$ (5)
+By computing contextual gain for both the original context and for the expectation over random contexts we can quantify how much more important the original context is for AOP-$\Delta$.
+We plot the results for this experiment in Figure 6B.
+The true context improves isolated AOP-$\Delta$ in 65.1% of the cases, whereas random contexts lead to an improvement for only 39.6% (for 33.2% only the true context leads to an improvement whereas random contexts don’t).
+Figure 8: Four sentences annotated with the Integrated Gradient attribution scores, that each denote a specific different way in which part-of-speech plays a role in contextual adjective order.
+Context Properties
+In Figure 7 we plot the impact of context length against AOP-% and Context Gain.
+Longer contexts lead to better performance as a whole (7A), and lead to larger differences between the contextual and isolated AOP scores (7C).
+This effect gradually increases with model size: longer contexts have the greatest effect for the largest model.
+One way that contextual cues may help a model is to disambiguate adjective orders with high entropy.
+We compute adjective order entropy based on the $\textsc{a}_{1}\textsc{a}_{2}$ bigram counts from Infini-gram (§5.1).
+In Figure 7B we demonstrate that model performance drops significantly for adjective pairs with high entropy: these are the hardest cases to get right.
+Surprisingly, however, the context gain does not increase for high entropy items (7D), indicating that the model can not leverage contextual cues for disambiguation.
+6.2 Uncovering Context Cues
+Thus far we have only examined the role of context on a macro-level.
+We now present an approach to identify specific words in the context that are of importance to the model’s predictions.
+For this we make use of feature attribution methods, which express the relative importance of input features to a prediction.
+Procedure
+Our procedure for identifying salient context cues is inspired by Sarti et al.
+(2024), who use feature attribution techniques to identify contextual cues in machine translation.
+Key to this approach is the usage of contrastive explanations (Yin and Neubig, 2022), where the contribution of a context token is expressed with respect to the difference of two possible outputs (in our case the two adjective orders).
+  1. 1.
+    Using the context gain scores of Eq.
+5, we create a subset of context-dependent constructions by taking the top 20% of items in CAP.
+These are the items for which the sentence context had the greatest impact on the model’s adjective order preference.
+  2. 2.
+    We compute the contribution of the tokens in the context using Integrated Gradients (Sundararajan et al., 2017) with respect to the AOP-$\Delta$ scores.
+[^2] Scores are obtained using inseq (Sarti et al., 2023), with 0-valued baselines.
+We sum over the embedding dimension to aggregate subword-level attributions, and sum subword attributions to obtain token-level attributions.
+Since Integrated Gradients is linear we can obtain contrastive explanations by subtracting the attribution scores of the swapped phrase from the scores of the original phrase.
+Positive contribution scores indicate that that token boosts the original order, whereas a negative score indicates a contribution to the opposite order.
+[^3] Feature attribution methods have been under extensive scrutiny in recent years: the complexity of large-scale models makes it challenging to capture their behaviour in a singular importance score, which will lead to disagreement between different methods (Neely et al., 2022).
+Integrated Gradients have nonetheless been shown to be reliable estimators of feature importance (Mariotti et al., 2024; Sarti et al., 2024), and therefore serve as a useful baseline for our experimental setup.
+We acknowledge though that different attribution methods may yield different insights.
+  3. 3.
+    We filter for cases where the top contributing token had a score of at least 1, indicating that there is a salient cue present in the context.
+  4. 4.
+    In order to map our results to a ‘global’ interpretation we POS tag the contexts using SpaCy’s en_core_web_lg tagger (Honnibal et al., 2020).
+This allows us to express the model behaviour in terms of POS classes, rather than on a lexical level.
+Figure 9: Distribution of the most important token in the context, grouped by POS tag.
+Each point indicates the number of times that POS tag was the most important token, and the median (relative) distance to the adjective phrase.
+POS tags with a frequency below 30 are filtered out.
+Results
+We perform our procedure on the Pythia-1.4b model.
+In Figure 8, we show four annotated sentences that are sampled randomly from the corpus.
+Each sentence exemplifies the way different part-of-speech classes can play a role in influencing contextual AOP.
+It can be seen that a context may contain ‘competing’ cues that prefer the natural and swapped adjective order, highlighting that context can impact model behaviour in different ways.
+To aggregate this result, we measure the amount of times a POS tag was the most important token in the context, and its distance from the adjective phrase.
+We visualize this in Figure 9.
+From this plot we conclude that there are two types of contextual cues that play a role:
+  1. 1.
 […]
-We compute the expectation over a random sample of 500 contexts, taken from CAP ($\mathcal{C}$).
-To express the relative impact of adding context, we regress out a model’s isolated AOP-$\Delta$ score by subtracting it from the contextual scores, which we call $c\textup{-}\Delta$:
-$c\textup{-}\Delta(\bullet|c)=\textup{AOP-}\Delta(\bullet|c)-\textup{AOP-}\Delta(\bullet)$ (5)
-By computing this score for both the original context and for the expectation over random contexts we can quantify how much more important the original context is for AOP-$\Delta$.
-We plot the results for this experiment in Figure 6B.
-The true context improves isolated AOP-$\Delta$ in 65.1% of the cases, whereas random contexts lead to an improvement for only 39.6%.
-Furthermore, for 33.2% the true context leads to an improvement whereas random contexts don’t: this subset can serve as a valuable starting point for investigating what properties of the context are important for adjective order.
-Conclusions
-AOP in Pythia is context-sensitive.
-Providing the model with the sentence where the adjective combination appears increases the model’s preference towards the target adjective order – both compared to the noun phrase in isolation and to the noun phrase in a random context.
-This result points in the direction of a more linguistically interesting mechanism for AOP than reliance on $n$-gram counts in training data alone.
-In particular, it shows a non-trivial interaction between the noun phrase and bigger context that turns out relevant for adjective placement.
-Establishing this opens up a possibility of studying the forms and types of this contextual interaction – a potential novel contribution to linguistic theories of adjective order.
-We leave a systematic discovery of types of AOP context-dependence open for future work, and focus on a small-scale qualitative study for now.
-An inspection of context-sensitive cases from CAP reveals at least two ways in which context matters for AOP:
-1. Structural parallelism
-The ‘canonical’ word order can be overridden when context contains a parallel noun phrase with adjectives in non-canonical order (the example shows contextual AOP according to Pythia-1.4b):
-  • a)
-    The green big cars AOP-$\Delta$ 4.2
-  • b)
-    The story is not about small red houses, it’s about big green cars AOP-$\Delta$ 6.63
-2. Property saliency
-The property salient in the context brings the adjective describing this property to the left edge of the noun phrase.
-If the sentence is about color, color adjective will come first; if the sentence is about size, it will bring size adjectives to the left (again, orders preferred by Pythia-1.4b):
-  • a)
-    They asked my favourite color of houses and I said I liked red small houses AOP-$\Delta$ 0.32
-  • b)
-    They asked my favourite size of houses and I said I liked small red houses AOP-$\Delta$ 1.61
-These observations, of course, barely scratch the surface of possible contextual AOP interactions, but exemplify the power of LMs in incorporating contextual information into word order judgments.
+  2. 2.
+    Semantic cues: Non-immediate context can play a role in AOP under a different mechanism.
+Our results highlight the role of adjectives and nouns that occur quite early in the sentence.
+In the case of adjectives this is strongly driven by semantic similarity: top-scoring adjectives in the context have on average a cosine similarity of 0.15 with the first adjective in the phrase, and only 0.07 with the second adjective.
+[^4] Cosine similarity is computed using the input embeddings of the LM, summing subword embeddings where necessary.
+6.3 Conclusions
+The results in this section show that adjective order behaviour in LMs is context-sensitive.
+Building on the aggregate effects reported earlier (Figure 3C), the attribution analysis identifies two classes of contextual cues that shape model preferences: local collocational cues in the immediate vicinity of the adjective phrase, and earlier semantic cues that bias adjective ordering based on semantic similarity.
+Importantly, these findings do not constitute an explanatory account of adjective order in human language.
+Rather, they illustrate how language models can serve as a tool for probing which aspects of sentence context are informative for adjective order preferences.
+In doing so, this analysis highlights contextual factors that are largely absent from existing theories of adjective order, and may serve as a starting point for future empirical studies on the role of context in human adjective order preferences.
 7 Discussion
-Abstract Notions of Adjective Order
-In this paper, we looked for AOP generalization in LMs.
-We were interested in whether abstract driving forces are at play in LM predictions beyond simple $n$-gram counts in training data.
-From our experiments we conclude that LMs show limited AOP generalization.
-Here, we want to highlight an important question that is rarely brought up in linguistic generalization discussions in NLP: How general and abstract is human linguistic behavior?
-Linguistic literature suggests that human language production, including word order preferences, is guided by a mix of memorization and abstract principles (O’Donnell, 2015).
-Speakers tend to recycle more routinized and therefore more accessible combinations (MacDonald, 2013), guided by sequence frequencies in linguistic input that, in turn, form sequential associations that lead to production automatizing (Bybee, 2010; Diessel, 2019).
-This is compatible with a picture of language acquisition where knowledge of grammar is formed based on encountered exemplars of usage, which are stored in memory (Bybee, 2010) and serve as basis of generalization (Goldberg, 2006), but can also serve as patterns for routinized language production.
+Generalization and Memorization in AOP
+In this paper, we investigated the extent to which language models generalize adjective order preferences beyond direct memorization of training data statistics.
+Our results suggest that while models exhibit some degree of generalization, this generalization is limited and coexists with a strong reliance on distributional information.
+Here, we want to highlight an important question that is rarely brought up in linguistic generalization discussions in NLP: How general and abstract is human linguistic behaviour?
+The linguistic literature suggests that human language production, including word order preferences, is guided by a mix of memorization and abstract principles (O’Donnell, 2015).
+Speakers tend to recycle more routinised and therefore more accessible combinations (MacDonald, 2013), guided by sequence frequencies in linguistic input that, in turn, form sequential associations that lead to production automatizing (Bybee, 2010; Diessel, 2019).
+This is compatible with a picture of language acquisition where knowledge of grammar is formed based on encountered exemplars of usage, which are stored in memory (Bybee, 2010) and serve as basis of generalization (Goldberg, 2006), but can also serve as patterns for routinised language production.
 So, human linguistic knowledge and language production is not totally abstract and is not based exclusively on generalization.
 This has been discussed across linguistic phenomena and has been suggested for word order (MacDonald, 2013; Levshina et al., 2023), although not for adjective order specifically.
-When investigating generalization in LMs it is therefore important to make a distinction between the degree to which they create generalizations for a particular phenomenon, and how similar this generalization is to human behavior.
-In principle, one could even find cases of LM overgeneralization, if we are interested in how human-like LMs are linguistically.
-Importantly, in order to answer the question of human-like generalization, we would need data on the trade-off in memorization and generalization in human linguistic behavior.
-Such data does not exist yet for adjective order preferences.
+When investigating generalization in LMs it is therefore important to make a distinction between the degree to which they create generalizations for a particular phenomenon, and how similar this generalization is to human behaviour.
+In principle, one could even find cases of LM overgeneralization.
+Importantly, in order to answer how human-like a model’s generalization behaviour is, we would need data on the trade-off in memorization and generalization in human linguistic behaviour: an empirical question for which detailed human data on adjective order preferences is currently lacking.
+Corpus Interventions
+Our analysis of AOP generalization is constrained by the fixed training corpus the models are trained on.
+More detailed insights into this question can be obtained by an intervention-based approach, in which we filter out specific constructions from the training data (Jumelet et al., 2021; Misra and Mahowald, 2024; Patil et al., 2024; Müller-Eberstein et al., 2025).
+An interesting experiment in this direction would be to remove all sentences containing double adjectives from the training corpus, and measure whether the remaining distribution of adjectives contains a sufficient signal to acquire adjective order constraints.
+Cross-lingual Extensions
+Our analysis is limited to adjective order preferences in English, which exhibits ordering patterns that differ from those found in other languages.
+Cross-linguistic research has documented both variation and regularities in adjective ordering, motivating general theories that go beyond any single language (Cinque, 2010; Culbertson et al., 2020).
+We leave a systematic cross-linguistic investigation of adjective order preferences in language models to future work.
 LMs and Linguistic Theory
@@
 There are various reasons for this.
-On the one hand, the cognitive plausibility of current LMs remains questionable, both in terms of the scale at which they are trained (Huebner et al., 2021; Warstadt et al., 2023) and their inductive biases that depend strongly on model architecture and are not reflective of human language processing (Wilson and Frank, 2023; Oh et al., 2024).
+On the one hand, the cognitive plausibility of current LMs remains questionable, both in terms of the scale at which they are trained (Huebner et al., 2021; Wilcox et al., 2025) and their inductive biases that depend strongly on model architecture and are not reflective of human language processing (Wilson and Frank, 2023; Oh et al., 2024; Padovani et al., 2025).
 On the other hand, their black box nature makes it highly challenging to extract linguistic patterns that can be tested against existing linguistic theories.
-While advancements in interpretability have made this latter point less problematic, the number of works explicitly connecting LM behavior to human language processing remains few and limited to simple syntactic phenomena (e.g., Lakretz et al., 2021).
-In this paper, we explore the possibility that LMs may ‘hide’ a theory of adjective order, by encoding features that are not covered by linguistic theories on adjective order.
-We show that LMs far exceed the predictive power of prior cognitive predictors, which raises the question whether this increase stems from meaningful interpretable features.
-While a lot of the LM predictions can be explained by collocational patterns—and those statistics alone do not explain the distribution of adjective orders—we do find some generalization to unseen adjective combinations and a linguistically meaningful pattern in the way that context boosts AOP.
-We hope our work can serve as an inspiration for using LMs as tools for measuring context-dependence, which is hard to quantify with traditional quantitative methods.
-Corpus Interventions
-Our analysis of AOP abstractness is constrained by the fixed training corpus the models are trained on.
-More detailed insights into this question can be obtained by a intervention-based approach, in which we filter out specific constructions from the training data (Jumelet et al., 2021; Misra and Mahowald, 2024; Patil et al., 2024).
-An interesting experiment in this direction would be to remove all sentences containing double adjectives from the training corpus, and measure whether the remaining distribution of adjectives contains a sufficient signal to acquire adjective order constraints.
-Limitations
-Our analysis only focuses on a single language model family (Pythia), in a a single language (English).
-To test the cross-lingual properties of adjective order, a wider range of language models should be examined.
-Our double adjective extraction method (§3.2) depends on the quality of the dependency parser that was used, and as such contains inaccuracies.
-Furthermore, a downside of using such a broad-coverage approach is that it becomes harder to draw targeted conclusions about the linguistic properties of cases on which the model fails or succeeds.
-Annotating cases containing properties that have been shown to be important for adjective order in the linguistics literature would provide deeper insights into model behavior on a linguistic level.
+In this paper, we treat language models not as explanatory theories of adjective order, but as empirical tools for analysing how soft word order constraints are acquired by LMs.
+Although language models substantially outperform individual cognitive predictors proposed in prior work, predictive accuracy alone does not constitute an explanation of adjective order preferences in human language.
+Instead, we view their value as lying in the combination of behavioural analysis with interpretability methods, which makes it possible to probe what sources of information influence model decisions (Futrell and Mahowald, 2025).
+In particular, the attribution-based analysis in § 6 illustrates how such methods can be used to identify contextual cues that systematically modulate adjective order preferences—factors that are largely absent from existing accounts and that may serve as targets for future empirical investigation in human language.
+Acknowledgements
+We gratefully thank the anonymous reviewers for their feedback on an earlier version of the paper, which greatly helped in shaping the story.
+We thank Kanishka Misra, Kyle Mahowald and Jennifer Culbertson for the inspiration that our adjacent collaboration has provided for finalizing this paper.
+Jaap Jumelet is supported by NWO grant VI.Vidi.221C.009.
 Appendix A Complementary Results
-[…]
-We provide a sample of CAP examples in Table 2, categorised by the contextual and isolated AOP scores of Pythia-12b.
-[…]
-CAP order always preferred
-1. In solution, lead(II) hydroxide is a somewhat weak base, forming lead(II) ion, Pb, under weakly acidic conditions. $16.5$ $18.8$
-2. It has the only deep water port in Cambodia. $16.9$ $15.4$
-3. The greatest social reforms in Denmark are certainly the work of the last half-century. $15.4$ $16.3$
-4. Icterids are unusual in songbirds because they have considerable sexual dimorphism. $16.9$ $14.5$
+A.1 CAP Examples
+We provide a sample of CAP examples in Table A.1, categorised by the contextual and isolated AOP scores of Pythia-12b.
+CAP Sentence AOP()($\bullet$) AOP(|c)($\bullet$|c)
+black       CAP order always preferred
+1. The seat comprises the following electoral wards $14.5$ $23.6$
+2. The local wildlife includes leopards, deer, monkeys, several kinds of pheasants, the marbled polecat and the increasingly rare flying squirrel $11.1$ $24.9$
+3. He was elected as leader of the Labour Party in the first round of voting using the single transferable vote $13.4$ $21.7$
+4. Hoggett and mutton can taste more flavorful than lamb because they contain a higher concentration of species- characteristic fatty acids $12.8$ $22.1$
+5. Jimmy seemed ready to take up that challenge, and entered upon an argument calculated to prove that he was a mild mannered individual $14.4$ $18.4$
 Swapped order always preferred
-5. Couches are usually bought in a set together with cushions, which give them a bouncey and decorative comfortable touch. $-7.1$ $-6.8$
-6. No, its 10 feet tall, and it’s a red big monster, demon like. $-7.3$ $-5.9$
-7. It does this by following the grammatical basic rules of syntax. $-5.8$ $-6.0$
-8. The daily average attendance from January through November 2010 was 22,133 people a day. $-5.8$ $-5.7$
+6. Mm, well, the extraction of foreign rectal bodies 11.3-11.3 11.3-11.3
+7. did they sing it’s an old grand flag 7.7-7.7 14.1-14.1
+8. His great jaw protruded frightfully, with the upper thin lip 6.6-6.6 6.4-6.4
+9. And Stuart sat down in, because they got these wooden old chairs 6.9-6.9 6.0-6.0
+10. It does this by following the grammatical basic rules 5.9-5.9 6.1-6.1
 Context improves AOP
-9. Being a spherical 3-manifold, it is the only homology 3-sphere, besides the 3-sphere itself, with a finite fundamental group. $0.9$ $16.8$
-10. Since the 2015 Styria municipal structural reform, it is part of the municipality Birkfeld. $-0.5$ $15.3$
-11. I won it at a fair last night. $0.2$ $15.4$
-12. I say that the man who does not play whist lays up a sad old age for himself. $2.6$ $17.5$
+11. Be it thy course to busy giddy minds $1.9$ $26.2$
+12. I have not passed through fire and death to bandy crooked words $5.3$ $29.2$
+13.
+"I wish to preach, not the doctrine of ignoble ease, but the doctrine of the strenuous life; the life of toil and effort; of labor and strife; to preach that highest form of success which comes, not to the man who desires mere easy peace $2.3$ $21.8$
+14. The "Yonsei" are the subject of on- going academic research 0.2-0.2 $18.1$
+15. The city flag of Portland, Oregon, consists of a green field on which is placed a white 4-pointed star from which radiate blue stripes, each bordered by L- shaped yellow lines 1.0-1.0 $16.9$
 Context worsens AOP
-13. This is an ability which may have been present in their last common ancestor in the Archaean. $12.3$ $4.0$
-14. Jacob, you’ve had a complex partial seizure, which can cause psychosis, including religious psychosis. $7.3$ $0.3$
-15. He was also a right-arm medium fast bowler with three wickets in test matches to his credit. $6.2$ $-0.6$
-16. He wheeled round sharply, and distinguished her lying with helpless outspread arms on the couch. $4.0$ $-2.8$
+16. One expected that it earned a alone decisive battle $9.1$ 1.1-1.1
+17. Meager is known for its last major eruption $12.8$ $4.8$
+18. As a matter of fact, no body in the universe revolves about the exact geometrical center $7.6$ 0.4-0.4
+19. E major is a musical major scale $4.8$ 2.8-2.8
+20. That worthy was seated before a table spread with papers, and as Tom entered or rather was pushed into his presence he compressed his beetling black brows $5.1$ 2.5-2.5
 Table 2: Most salient examples of the four quadrants in Figure 6.
-A.2 Correlations to Cognitive Predictors
-Correlations during training of AOP-$\Delta$ scores (Pythia-1.4b), with scores derived from various cognitive predictors.
+A.2 Overlap and Correlations to Cognitive Predictors
+Overlap between various cognitive predictors and AOP-$\Delta$ scores of Pythia-1.4b::
+[Uncaptioned image][Uncaptioned image]
+Correlations during training of AOP-$\Delta$ scores (Pythia-1.4b), with scores derived from various cognitive predictors: [Uncaptioned image]
 A.3 Relative Number of Bigram Occurrences
@@
 The dotted line denotes the average AOP-$\Delta$ across all CAP items.
+A.6 Context Length Impact
+Relation between context length and contextual AOP-$\Delta$ scores, averaged over randomly sampled contexts.
+The context length effect shown in Figure 7 disappears when averaging over random contexts, showing that contextual gain is not driven by context length itself.